### v1

Abstract
Image restoration involves recovering a high-quality clean image from its degraded version, which is a fundamental task in computer vision. Recent progress in image restoration has demonstrated the effectiveness of learning models capable of addressing various degradations simultaneously, i.e., the All-in-One image restoration models. However, these existing methods typically utilize the same parameters facing images with different degradation types, which causes the model to be forced to trade off between degradation types, therefore impair the total performance. To solve this problem, we propose HAIR, a Hypernetworks-based plug-in-and-play method that dynamically generated parameters for the corresponding networks based on the contents of input images. HAIR consists of 2 main components: Classifier (Cl) and Hyper Selecting Net (HSN). To be more specific, the Classifier is a simple image classification network which is used to generate a Global Information Vector (GIV) that contains the degradation information of the input image; And the HSNs can be seen as a simple Fully-connected Neural Network that receive the GIV and output parameters for the corresponding modules. Extensive experiments shows that incorporating HAIR into the architectures can significantly improve the performance of different models on image restoration tasks at a low cost, although HAIR only generate parameters and haven’t change these models’ logical structures at all. With incorporating HAIR into the popular architecture Restormer, our method obtains superior or at least comparable performance to current state-of-the-art methods on a range of image restoration tasks. Code and pre-trained checkpoints are available $\underline{\textbf{Code and pre-trained checkpoints are available here.}}$
1 Introduction
Image restoration stands as a pivotal challenge within the field of computer vision, dedicated to the reconstruction of high-quality images from their degraded states. The presence of adverse conditions such as noise, haze, or rain can severely diminish the effectiveness of images for a variety of applications, such as autonomous navigation [1, 2], augmented reality [3, 4, 5]. Developing robust image restoration methods is thus an imperative pursuit. The utilization of deep learning into this domain has yielded remarkable progress, as evidenced by a suite of recent methodologies [6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Nonetheless, the predominant approach in current research is to employ task-specific models, each tailored to address a particular type of degradation [16, 7, 12, 13, 8]. This tailored approach, while precise, presents a constraint in terms of versatility, as it restricts the applicability of models to scenarios with varied or unknown degradations [17, 18, 19].
To overcome this limitation, the research community has increasingly turned its attention to the development of All-in-One image restoration models. These models are designed to adeptly manage a spectrum of degradations in a single framework. Pioneering efforts in this area [20, 21, 1, 22, 23, 24, 25, 26, 15] have utilized a variety of advanced techniques such as contrastive learning [20], meta-learning [25], visual prompting methods [24, 27, 28, 14]. These approaches have undoubtedly made substantial contributions to the field, propelling it forward with notable advancements.
Figure 1: The difference between our method and previous methods. (a) Previous All-in-One image restoration methods. For images with various degradations, these methods utilize a single model with the same parameters. (b) Our proposed HAIR. Given a certain degraded image, we first use a fixed Extraction module to extract the features, and feed the extracted feature into the Hypernetwork to generate the dynamic parameters for the Fusion module, and finally get the restored image by the generated parameters.
Despite the notable successes of these methods, they share a common drawback: they rely on a single model with fixed parameters to address a variety of degradations. This one-size-fits-all approach can hinder the model’s effectiveness when dealing with multiple degradations simultaneously. For instance, when viewed through the lens of frequency domain analysis, haze is characterized as low-frequency noise, in contrast to rain, which is considered high-frequency interference. An effective dehazing model acts as a low-pass filter, preserving high-frequency details, whereas deraining requires the opposite—enhancing the high-frequency components. Consequently, a model must strike a balance between these conflicting demands of different degradations, which may lead to suboptimal performance.
To mitigate the aforementioned issue, we propose HAIR in this paper. Our fundamental concept is the dynamic generation of parameters based on the degradation information present in the input image. HAIR employs Hypernetworks [29], a trainable neural network that outputs parameters for other neural networks, to take the degradation information from the input image and produce corresponding parameters. Specifically, for a given unknown degraded image, we first utilize a Classifier, similar to those used in image classification networks, to obtain its Global Information Vector. This vector is then passed to the Hyper Selecting Net, which is a Fully Connected Network, to generate the requisite parameters, as illustrated in Fig. 1. Thereby, our method can dynamically adapt to various degradations. With these dynamically-parameterized modules, we ultimately achieve the restored image. In brief, our contributions include:
  • We propose HAIR, a novel Hypernet-based blind All-in-One image restoration method capable of dynamically generating parameters based on the input image’s characteristics. By incorporating HAIR with Restormer [12], our method achieves performance that is superior or at least comparable to that of current state-of-the-art methods across a variety of image restoration tasks.
  • Our proposed HAIR, which comprises a Classifier and Hyper Selecting Nets, functions as a plug-in-and-play module. It can be effortlessly integrated into any existing restoration network without altering any logical structures, thereby enhancing the network’s generalization and robustness at a minimal additional cost. Extensive experiments indicate that HAIR possesses strong universality.
  • To the best of our knowledge, our method is the first to employ data-adaptive Hypernetworks in image restoration tasks, demonstrating their strong potential and offering a novel perspective for advancing image restoration capabilities.
Figure 2: The figure presents tSNE plots of the degradation embeddings utilized in PromptIR [24] and our HAIR. Each distinct color represents a unique degradation type. It becomes evident that HAIR excels not only in recognizing various degradation types, such as noise, rain, and haze, but also in distinguishing between the same type of degradation at varying intensities, exemplified by noise with standard deviations of [15, 25, 50]. This capability underscores HAIR’s remarkable adaptability and generalization when confronted with diverse degradation scenarios.
2 Related Works
2.1 All-in-One Image Restoration
While single degradation methods do achieve great success [9, 10, 11, 7, 12, 13], blind all-in-one image restoration, which aims to utilize a single deep blind restoration model to tackle multiple degradation types simultaneously, gaining more attention recently  [1, 2, 30, 24, 25, 26]. The seminal work AirNet [20] achieves blind All-in-One image restoration by employing contrastive learning to extract degradation representations from corrupted images, which are subsequently utilized to restore the clean image. Next, IDR [25] decomposes image degradations into their underlying physical principles, achieving All-in-One image restoration through a two-stage process based on meta-learning. With the rise of LLMs [31], Prompt-based learning [24, 27, 14] also emerged as a promising direction in computer vision tasks. Despite the success these methods achieved, they typically use the same parameters for distinct degradations and inject the degradation information as conditional embeddings into the model, which can impair the performance of the models. In contrast, our proposed HAIR can adaptively generate specific parameters for the given degraded image, resulting in better generalization to unknown degradations. [32] explains in theory why using Data-conditional Hypernetworks to generate adaptive parameters work better than injecting conditional embeddings into the model.
2.2 Image Classification
Image Classification is a fundamental task in Computer Vision, with the goal of categorizing images based on their content [33]. Typically, the process of Image Classification involves two key steps: feature extraction and the subsequent classification itself. The primary challenge in this task is feature extraction, which involves converting an image into a vector that encapsulates its most critical information. While traditional methods relied on handcrafted techniques [34], the advent of neural networks has revolutionized feature extraction [35, 36, 37, 38]. Given that these vectors are capable of classifying hundreds of categories [39], they must contain the essential information of the images. This concept extends to the classification of degraded images as well. For instance, if we categorize degraded images into types such as noisy, rainy, or hazy, and by levels of degradation like slightly noisy, moderately noisy, and very noisy, the same vectors should be effective in classifying these variations. The rationale is that recognizing different degradations is generally less complex than identifying hundreds of distinct objects. Therefore, our proposed method employs a Classifier to obtain a Global Information Vector (GIV). This GIV, rich in the global information of the image, including its degradation characteristics, is then utilized to generate parameters tailored to the specific degradation present in the image.
2.3 Data-conditioned Hypernetworks
Hypernetworks [29] are a class of neural networks designed to generate weight values (parameters) for other networks. They can be broadly classified into three main types: Task-conditioned, Data-conditioned, and Noise-conditioned hypernetworks [40]. Of these, Data-conditioned hypernetworks are particularly noteworthy for their ability to generate weights contingent upon the distinctive features of the input data. This capability allows the neural network to dynamically adapt its behavior in response to specific input patterns or characteristics, fostering flexibility and adaptability within the model. Consequently, this results in enhanced generalization and robustness. Data-conditioned Hypernetworks have found applications across the field of computer vision, including semantic segmentation [41] and image editing [42]. Despite previous attempts to integrate Hypernetworks into Image Restoration [43], these have primarily leveraged Task-conditioned hypernetworks. Such an approach can’t fully harness the potential of hypernetworks for the task of image restoration. To the best of our knowledge, this work pioneers the introduction of Data-conditioned hypernetworks to the domain of image restoration, unlocking new possibilities for the field.
Figure 3: The overall framework of HAIR. In this paper, we incorporate HAIR into Restormer [12], which has demonstrated robust performance in all-in-one restoration scenarios [24, 26, 28]. HAIR introduces a Classifier that yields a Global Information Vector (GIV) $\mathbf{V_{g}}$ from the high-level feature $\mathbf{F_{g}}$ encapsulating the input image’s degradation information. Additionally, the Hyper Selecting Nets (HSNs) dynamically generate weights for the Transformer Blocks based on GIV, customizing the restoration process to address the specific degradation characteristics of each image. This innovative approach allows models to adapt to various image restoration tasks without altering the original architecture.
3 Method
In this section, we outline the core principles of our proposed effective all-in-one image restoration method.
Overall Pipeline As shown in Fig. 3, our network architecture is consistent with Restormer [12]. Given a certain degraded image input $\mathbf{X}\in\mathbb{R}^{H\times W\times 3}$, HAIR utilizes a $3\times 3$ convolution to transform $\mathbf{X}$ into feature embeddings $\mathbf{F_{0}}\in\mathbb{R}^{H\times W\times C}$, where $C$ denotes the number of channels. These feature embeddings then proceed through a 4-level hierarchical encoder-decoder, resulting in deep features $\mathbf{F_{r}}$. Each level of the encoder-decoder incorporates several Transformer blocks, with an increasing number from the top to the bottom level, ensuring computational efficiency. The left three blue "Transformer Blocks" function as an Encoder, designed to extract features from $\mathbf{F_{0}}$ and ultimately produce a global feature map $\mathbf{F_{g}}\in\mathbb{R}^{\frac{W}{8}\times\frac{H}{8}\times 8C}$ with a large receptive field. $\mathbf{F_{g}}$ is then input into a Classifier to yield a Global Information Vector (GIV) $\mathbf{V_{g}}\in\mathbb{R}^{2C}$. The right four orange "HyperTrans Blocks" operate as a Decoder, aiming to adaptively fuse features at each Decoder level based on degradation, culminating in a restored image $\hat{\mathbf{X}}$. Specifically, at each Decoder level, the HyperTrans Block accepts two inputs: the input feature and the GIV $\mathbf{V_{g}}$. $\mathbf{V_{g}}$ is fed into the Hyper Selecting Net to generate weights for the corresponding Transformer Blocks, which are then applied to the input feature to produce the output. It is important to note that all HyperTrans Blocks use the same $\mathbf{V_{g}}$ derived from $\mathbf{F_{g}}$. Since the weights in the Decoder are generated based on $\mathbf{V_{g}}$, which contains the degradation information of $\mathbf{X}$, our method exhibits strong generalization and adaptability, leading to excellent restoration performance.
3.1 Classifier
As outlined in Section 2.2, our approach to extracting degradation information involves designing a straightforward Classifier akin to those used in image classification tasks. We commence by inputting $\mathbf{F_{g}}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 8C}$ into the backbone, which is comprised of multiple ResNet Blocks [35] followed by ×$\times 2$ Downsampling. This process is aimed at downsizing the spatial resolution while distilling the essential information, as depicted by the following equations:
$\mathbf{F_{g}^{1}}$ $=\text{Downsampling}(\text{ResNetBlock}(\mathbf{F_{g}})),$ (1)
$\mathbf{F_{g}^{t}}$ $=\text{Downsampling}(\text{ResNetBlock}(\mathbf{F_{g}^{t-1}}))),\quad\text{for}\ t=1,2,3.$
After three iterations, we obtain $\mathbf{F_{g}^{3}}\in\mathbb{R}^{\frac{H}{64}\times\frac{W}{64}\times 2C}$, which serves as the input for generating the GIV $\mathbf{V_{g}}\in\mathbb{R}^{2C}$:
$\mathbf{V_{g}}=\text{GAP}(\mathbf{F_{g}^{3}}),$ (2)
where GAP denotes Global Average Pooling. The inclusion of GAP ensures that $\mathbf{V_{g}}$ remains a one-dimensional vector of consistent size, irrespective of the input image’s resolution. Unlike traditional image classification networks, there is no requirement for a Softmax layer here, as its application would confine the values of $\mathbf{V_{g}}$ within the [0,1] range, potentially restricting the parameter generation process. Moreover, the backbone can be substituted with other straightforward image classification networks such as VGG [44] and Inception [45], without any detrimental impact on performance, given that the classification of degradation types is a relatively simple task.
3.2 Hyper Selecting Net
Upon obtaining the Global Information Vector $\mathbf{V_{g}}$ from the Classifier, it is fed into our Data-conditioned Hypernetworks, known as Hyper Selecting Nets, to generate weights for the corresponding Transformer Blocks. Given $\mathbf{V_{g}}\in\mathbb{R}^{2C}$, a Selecting Vector (SV) $\mathbf{V_{s}}\in\mathbb{R}^{N}$ is initially computed as follows:
$\mathbf{V_{s}}=\sigma(\mathbf{FCNN}(\mathbf{V_{g}}))$ (3)
Here, $\sigma$ stands for a Softmax operation and $\mathbf{FCNN}$ denotes a simple Fully-Connected Neural Network. Subsequently, the parameters $\mathbf{w}$ are derived:
$\mathbf{w}=\sum_{i=1}^{N}\mathbf{V_{s}}^{i}\mathbf{W}_{i}$ (4)
In this formula, $\mathbf{V_{s}}^{i}$ represents the $i$-th element of $\mathbf{V_{s}}$. The matrix $\mathbf{W}\in\mathbb{R}^{N\times P}$, referred to as the Weights Box, comprises $\mathbf{W}_{i}\in\mathbb{R}^{P}$ as its $i$-th column. The hyperparameter $N$ influences the total number of parameters, with $P$ being the count of parameters required for one Transformer Block. With $\mathbf{w}$ determined, it is then applied as parameters for the Transformer Block:
$\mathbf{x}^{\prime}=\mathbf{Transformer\_Block}(\mathbf{x};\mathbf{w})$ (5)
Here, $\mathbf{x}$ signifies the input, and $\mathbf{x}^{\prime}$ is the resulting output. As the Transformer Blocks [12] are based on convolution, $\mathbf{w}\in\mathbb{R}^{P}$ is reshaped into 4-dimensional tensors to serve as convolution kernels for the Transformer Blocks. To economize on parameters, Transformer Blocks at the same decoder level share one Weights Box, and each Transformer Block is independently equipped with its own FCNN.
3.3 Rethinking Weights Generation via a Selecting Perspective
If we consider the Weights Box as part of the FCNN, it may not be immediately clear why such a direct derivation of weights from a neural network works well. This seeming lack of intuitiveness is what we aim to address. Before delving into the explanation, we establish a foundational proposition:
Proposition 1
Convolution operations exhibit the distributive law over addition. Let $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$ be the input feature, and let $w_{i},\ i=1,2,\cdots,n$ represent the convolution kernels. The law is mathematically expressed as:
$\mathbf{x}*(\sum_{i=1}^{n}w_{i})=\sum_{i=1}^{n}(\mathbf{x}*w_{i})$ (6)
where ’*’ denotes the standard 2-dimensional convolution.
With this proposition, the following becomes evident:
$\mathbf{x}*\mathbf{w}$ $=\mathbf{x}*\left(\sum_{i=1}^{N}\mathbf{V_{s}}^{i}\mathbf{W}_{i}\right)$ (7)
$=\sum_{i=1}^{N}\left(\mathbf{x}*(\mathbf{V_{s}}^{i}\mathbf{W}_{i})\right)$
$=\sum_{i=1}^{N}\mathbf{V_{s}}^{i}\left(\mathbf{x}*\mathbf{W}_{i}\right)$
With the constraint $\sum_{i}^{N}\mathbf{V_{s}}^{i}=1$ due to the property of Softmax operation. In essence, the convolution with the generated weights is equivalent to a weighted sum of convolutions between the input $\mathbf{x}$ and each kernel $\mathbf{W}_{i}$ in the Weights Box. This process can be seen as a process of selecting convolution kernels. For instance, with a Weights Box $\mathbf{W}\in\mathbb{R}^{2\times P}$ that includes a low-pass $\mathbf{W}_{1}$ and a high-pass $\mathbf{W}_{2}$, the rainy input $\mathbf{x}_{1}$ would ideally have a larger $\mathbf{V_{s}}^{1}$ to filter out high-frequency noise, while a smaller $\mathbf{V_{s}}^{2}$ would help retain details. Conversely, for a hazy input $\mathbf{x}_{2}$, a larger $\mathbf{V_{s}}^{2}$ would be necessary to mitigate low-frequency haze. Although real-world degradations can be more complex, our Hyper Selecting Net is adept at selecting the appropriate weights. This insight into the selection process helps us understand HAIR: for each input, the Hyper Selecting Net produces a tailored Selecting Vector and adaptively chooses the most suitable convolution kernel. Given that core operations such as convolution and matrix multiplication follow the distributive law over addition, HAIR is versatile and can be integrated into more various architectures like Transformer [46] and Mamba [47].
3.4 A Brief Analysis of Model Complexity
In Section 2.3, we claim that our Hypernetworks-based method can work better than conditional embedding based methods like AirNet [20] and PromptIR [24], so we will give a simple proof for this. In the context of all-in-one image restoration, we aim to find a function $f:\mathbb{R}^{3HW}\xrightarrow{}\mathbb{R}^{3HW}$ to map a degraded image $\mathbf{X}$ to a restored image $\hat{\mathbf{X}}$, minimizing the distance between $\hat{\mathbf{X}}$ and the ground truth $\mathbf{Y}=y(\mathbf{X})$. (For the sake of discussion, the following tensors are generally regarded as flattened one-dimensional vectors) $(a)$ Mainstream methods today typically utilize a neural network $g(\mathbf{X},\mathbf{I(\mathbf{X})})$ to learn the mapping, where $\mathbf{I(\mathbf{X})}\in\mathbb{R}^{m}$ stands for the degradation information (degradation embedding) from the input image, such as the prompt in PromptIR [24]. These methods directly send $\mathbf{I}(\mathbf{X})$ together with $\mathbf{X}$ into the network to get $\hat{\mathbf{X}}$. $(b)$ For our HAIR, the neural network can be formulated as $h(\mathbf{X};\theta(\mathbf{I(\mathbf{X})}))$, where $\theta$ is a function that maps $\mathbf{I(\mathbf{X})}$ into the parameters of $h$. We define the distance between 2 functions as follows:
$d(g,y)=\min_{g}\max_{\mathbf{X}}\|g(\mathbf{X},\mathbf{I}(\mathbf{X}))-y(\mathbf{X})\|_{\infty}$ (8)
Given that vector functions can be complex, we define scalar function $f^{i}:\mathbb{R}^{M}\xrightarrow{}\mathbb{R},i=1,2,3\cdots M.$ of a vector function $f:\mathbb{R}^{M}\xrightarrow{}\mathbb{R}^{M}$. $i$ means $f^{i}$ produce the $i$-th value of $f$, then we have:
$d(g,y)$ $=\min_{g}\max_{\mathbf{X}}\max_{i}|g^{i}(\mathbf{X})-y^{i}(\mathbf{X})|$ (9)
$=\min_{g}\max_{i}\max_{\mathbf{X}}|g^{i}(\mathbf{X})-y^{i}(\mathbf{X})|$
$=\min_{g}\max_{\mathbf{X}}|g^{k(g)}(\mathbf{X})-y^{k(g)}(\mathbf{X})|$
Since the value range of $i$ is limited, given the function $g$, we can always find an integer $k(g)$ to replace $\max_{i}$. To complete the proof, we need some assumptions.
Assumption 1
The target function $y^{i}\in\mathcal{W}_{r,3HW},i=1,2,\cdots 3HW$. The Sobolev space $\mathcal{W}_{r,3HW}$ is the set of functions that are $r$-times differentiable with all $r$-th order derivatives being continuous and bounded by the Sobolev norm $\leq 1$, defined on $\mathbb{R}^{3HW}$.
Intuitively, since $y^{i}$ maps the degraded image to a single pixel of the clean image, a little disturbance on the degraded image should only bring a little difference to the single pixel, so $y^{i}$ should be smooth and differentiable or at least continue. What’s more, the domain and range of $y$ are restricted to $[0,1]^{3HW}$, with most of the pixels far more less than 1, the Sobolev norm of $y_{i}$ can be generally less than 1. So this assumption generally holds.
Assumption 2
For various functions (especially neural networks) $g$, the integer $k(g)$ in Eq. 9 remains the same.
In our context, the various $g$ actually stands for the same network with different weights during training. No matter how the parameters updates, the most "difficult" degraded image genrally remains the same one. (e.g., the one with very severe degradation) And within this "difficult" image, the most "difficult" pixel should also remains the same. (e.g., the pixel that varies dramatically from its clean version.) In this way, we simplify the vector function $g$ to scalar function $g^{k(g)}$. With these 2 assumptions, incorporating the Theorem 2, Theorem 3 Theorem 4 and all the assumptions in [32], we have the following conclusions:
Theorem 1
For conditional embedding based all-in-one image restoration methods $g(\mathbf{X},\mathbf{I(\mathbf{X})})$ like PromptIR [24], to achieve error $d(g,y)\leq\epsilon$, the complexity (number of parameters) of the corresponding model is at least: $N_{g}=\Omega(\epsilon^{-3HW-m})$
Theorem 2
For Hypernetworks based methods $h(\mathbf{X};\theta(\mathbf{I}(\mathbf{X})))$ like our HAIR, to achieve error $d(h,y)\leq\epsilon$, the complexity of the corresponding model is: $N_{\theta}=O(\epsilon^{-3HW/r})$
See paper [32] for detailed premises and proof of Theorem 1, 2. In this way, we show that to achieve the same error, HAIR will cost less parameters than conditional embedding based methods.
4 Experiments
We conduct experiments by rigorously adhering to the methodologies established in prior research on general image restoration [25, 24], under two distinct experimental configurations: (a) All-in-One and (b) Single-task. Under the All-in-One configuration, a singular model is trained to handle multiple degradation types, encompassing both three and five unique degradations. In contrast, the Single-task configuration involves the training of individual models dedicated to specific restoration tasks.
4.1 Experimenting Preparation
Implementation Details. Our HAIR framework is designed to be end-to-end trainable, eliminating the need for pre-training any of its components. The architecture comprises a 4-level encoder-decoder structure, with each level equipped with a varying number of Transformer blocks. Specifically, the number of blocks increases progressively from level-1 to level-4, following the sequence [4, 6, 6, 8]. HyperTrans Blocks are employed throughout level-4 and the decoding stages of levels 1-3. Additionally, the Weights Box parameter $N$ is set according to the sequence [5, 7, 7, 9] for each respective level. In the all-in-one setting, the model is trained with a batch size of 32, while in the single-task setting, a batch size of 8 is used. The network’s optimization is guided by an L1 loss function, employing the AdamW optimizer [48] with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate is set to $42e-$ 4 for the initial 150 epochs, and we change it to $52e-$ 5 for the final 10 epochs. To enhance the training data, input patches of size 128 x 128 are utilized, with random horizontal and vertical flips applied to the images to augment the dataset.
Datasets. We follow the approaches in previous studies [20, 24, 25] for our All-in-One and single-task experiments, using the following datasets: For single-task image denoising, we use the BSD400 [49] and WED [50] datasets, adding Gaussian noise at levels $\sigma\in\{15,25,50\}$ to generate training images. Testing is performed on the BSD68 [51] and Urban100 [52] datasets. For deraining, we employ the Rain100L dataset [53]. Dehazing experiments utilize the SOTS dataset [54]. Deblurring and low-light enhancement tasks use the GoPro [55] and LOL-v1 [56] datasets respectively. To create a comprehensive model for all tasks, we combine these datasets and train in a setting that covers three or five types of degradations. For single-task models, training is conducted for the same duration on the respective dataset.
Table 1: Comparison to state-of-the-art on three degradations. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the full RGB images. Best and second best performances are highlighted. Our method sets a new state-of-the-art on average across all benchmarks while being significantly more efficient than prior work. ‘-’ represents unreported results.
  Method Params. Dehazing Deraining Denoising Average
[…]
BRDNet [57] - 23.23 .895 27.42 .895 32.26 .898 29.76 .836 26.34 .693 27.80 .843
LPNet [58] - 20.84 .828 24.88 .784 26.47 .778 24.77 .748 21.26 .552 23.64 .738
FDGAN [59] - 24.71 .929 29.89 .933 30.25 .910 28.81 .868 26.43 .776 28.02 .883
DL [23] 2M 26.92 .931 32.62 .931 33.05 .914 30.41 .861 26.90 .740 29.98 .876
MPRNet [21] 16M 25.28 .955 33.57 .954 33.54 .927 30.89 .880 27.56 .779 30.17 .899
Restormer [12] 26M 29.92 .970 35.56 .969 33.86 .933 31.20 .888 27.90 .794 31.68 .910
AirNet [20] 9M 27.94 .962 34.90 .967 33.92 .933 31.26 .888 28.00 .797 31.20 .910
PromptIR [24] 36M 30.58 .974 36.37 .972 33.98 .933 31.31 .888 28.06 .799 32.06 .913
DaAIR[26] 6M 32.30 .981 37.10 .978 33.92 .930 31.26 .884 28.00 .792 32.51 .913
HAIR (ours) 29M 30.98 .979 38.59 .983 34.16 .935 31.51 .892 28.24 .803 32.70 .919
4.2 Results
All-in-One: Three Degradations. We conducted a comparative evaluation of our All-in-One image restoration framework against several specialized state-of-the-art methods, including BRDNet [57], LPNet [58], FDGAN [59], DL [23], MPRNet [21], AirNet [20], PromptIR [24] and DaAIR [26]. Each of these methods was trained to handle the tasks of dehazing, deraining, and denoising. As illustrated in Table 1, our approach excels as the leading contender, offering a significant average improvement of $0.64$ dB over PromptIR [24]. Our method notably outperforms existing benchmarks on BSD68 and Rain100L datasets, setting a new precedent by exceeding the performance of the previous best methods by margins of $0.19$ dB and $1.49$ dB respectively.
All-in-One: Five Degradations. Following recent studies [25, 26], we have extended our approach to include deblurring and low-light image enhancement, thereby broadening our three-degradation framework to a more complex All-in-One setting. As demonstrated in Table 2, our method excels by learning specialized models for each degradation type while effectively capturing the shared characteristics between tasks. It achieves an average improvement of $0.6$ dB over PromptIR [24] and $0.13$ dB over DaAIR [26], establishing a new benchmark for state-of-the-art (SOTA) performance across all five benchmarks. Additionally, we compared our method with general image restoration models trained in the same All-in-One setting. Notably, our method significantly outperforms our baseline, Restormer [12], in various tasks. This comparison highlights the strong capability of our method as a versatile plug-in-and-play module, enhancing the performance of existing restoration architectures with minimal integration complexity.
Table 2: Comparison to state-of-the-art on five degradations. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the full RGB images with $(^{\ast})$ denoting general image restorers, others are specialized all-in-one approaches. Best and second best performances are highlighted.
  Method Params. Dehazing Deraining Denoising Deblurring Low-Light Average
[…]
NAFNet [8] 17M 25.23 .939 35.56 .967 31.02 .883 26.53 .808 20.49 .809 27.76 .881
DGUNet [60] 17M 24.78 .940 36.62 .971 31.10 .883 27.25 .837 21.87 .823 28.32 .891
SwinIR [7] 1M 21.50 .891 30.78 .923 30.59 .868 24.52 .773 17.81 .723 25.04 .835
Restormer [12] 26M 24.09 .927 34.81 .962 31.49 .884 27.22 .829 20.41 .806 27.60 .881
DL [23] 2M 20.54 .826 21.96 .762 23.09 .745 19.86 .672 19.83 .712 21.05 .743
Transweather [1] 38M 21.32 .885 29.43 .905 29.00 .841 25.12 .757 21.21 .792 25.22 .836
TAPE [22] 1M 22.16 .861 29.67 .904 30.18 .855 24.47 .763 18.97 .621 25.09 .801
AirNet [20] 9M 21.04 .884 32.98 .951 30.91 .882 24.35 .781 18.18 .735 25.49 .847
IDR [25] 15M 25.24 .943 35.63 .965 31.60 .887 27.87 .846 21.34 .826 28.34 .893
PromptIR [24] 36M 30.61 .974 36.17 .973 31.25 .887 27.93 .851 22.89 .842 29.77 .905
DaAIR [26] 6M 31.97 .980 36.28 .975 31.07 .878 29.51 .890 22.38 .825 30.24 .910
HAIR (ours) 29M 30.62 .978 38.11 .981 31.49 .891 28.52 .874 23.12 .847 30.37 .914
Single-Degradation Results. To evaluate the effectiveness of our proposed method, we provide results in Table 3, which show the performance of individual instances of our method trained under a single degradation protocol. Specifically, the single-task variant dedicated to deraining consistently achieves higher performance than PromptIR [24] and DaAIR by margins of $1.96$ dB and $1.22$ dB, respectively. When applied to image denoising, our method also demonstrates superiority over the aforementioned approaches, with an improvement of $0.21$ dB and $0.47$ dB, respectively. These results underscore the significant enhancements in performance delivered by our method.
Table 3: Comparison to state-of-the-art for single degradations. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the full RGB images. Best and second best performances are highlighted. Our method excels prior work on deraining and denoising.
(a) Dehazing
Method SOTS
DehazeNet  [61] 22.46 .851
MSCNN  [62] 22.06 .908
EPDN  [63] 22.57 .863
FDGAN [59] 23.15 .921
Restormer [12] 30.87 .969
AirNet [20] 23.18 .900
PromptIR [24] 31.31 .973
DaAIR [26] 31.99 .981
HAIR (ours) 31.68 .980
(b) Deraining
Method Rain100L
DIDMDN  [64] 23.79 .773
UMR  [65] 32.39 .921
SIRR  [66] 32.37 .926
MSPFN  [67] 33.50 .948
Restormer [12] 36.74 .978
AirNet [20] 34.90 .977
PromptIR [24] 37.04 .979
DaAIR (ours) 37.78 .982
HAIR (ours) 39.00 .985
(c) Denoising
Method $\sigma$=15 $\sigma$=25 $\sigma$=50
CBM3D [68] 33.93 .941 31.36 .909 27.93 .833
DnCNN [69] 32.98 .931 30.81 .902 27.59 .833
IRCNN [6] 27.59 .833 31.20 .909 27.70 .840
FFDNet  [70] 33.83 .942 31.40 .912 28.05 .848
BRDNet [57] 34.42 .946 31.99 .919 28.56 .858
AirNet [20] 34.40 .949 32.10 .924 28.88 .871
PromptIR [24] 34.77 .952 32.49 .929 29.39 .881
DaAIR [26] 34.55 .949 32.24 .924 29.09 .872
HAIR (ours) 34.93 .953 32.70 .931 29.65 .885
Visual Results The visual results captured under three degradation scenarios are presented in Fig. 4. These results clearly demonstrate the superiority of our method in terms of visual quality: In the denoising task at $\sigma=25$, HAIR retains more fine details compared to other methods; in the deraining task, HAIR effectively removes all rain streaks, whereas the compared methods exhibit noticeable remnants; in the dehazing task, HAIR may not closely resemble the Ground Truth, yet it provides the most visually pleasing outcome and even clears the original haze present in the Ground Truth. Collectively, these observations confirm the efficacy of our approach in enhancing image quality across different degradation conditions.
Figure 4: Visual comparison of all-in-one methods in all-in-one degradation settings.
4.3 Ablation Studies
We have done several ablation studies to show the contributions and properties of our proposed HAIR.
Effectiveness of Classifier and Hyper Selecting Net (HSN) As detailed in Table 4, we study the impact of the proposed modules in four settings: (a) Here, the model aligns with the Restormer architecture [12]. (b) In this configuration, the Global Information Vector (GIV) is directly used as conditional embeddings for the Transformer Blocks in the Decoder, rather than for weight generation. (c) The GIVs are designated as independently trainable parameters, each randomly initialized, with the Decoder’s Transformer Blocks having their own distinct GIVs. (d) This setup incorporates both components. The results demonstrate the indispensability of both components. With the addition of only 3M parameters and no change to the logical structure, HAIR outperforms Restormer by 1.7 dB in PSNR, underscoring the simplicity and effectiveness of our proposed approach.
Position of Classifier (and HyperTrans Blocks) In our design approach, the Classifier is positioned after the first three Encoder levels, with the subsequent four Decoder levels employing HyperTrans Blocks, constituting a 3+4 configuration. The rationale for this design choice is demonstrated in Table 5. It is evident that the optimal placement for the Classifier is at the network’s midpoint. This positioning takes advantage of the expansive receptive field of the features post the network’s halfway point, while allowing the HSN to dynamically generate parameters for the remaining modules. Furthermore, it is crucial to strike a balance between the receptive field size of the features fed into the Classifier and the count of HyperTrans Blocks utilized, ensuring the model’s efficiency and adaptability.
Table 4: Impact of key components. Results are from single deraining task on Rain100L.
Setting Prams. Classifier HSN PSNR SSIM
(a) (baseline) 26M - - 36.74 .978
(b) 27M - 36.88 .979
(c) 28M - 36.76 .979
(d) (ours) 29M 39.00 .985
Table 5: Impact of position. Results are from Denoise task on Urban100 ($\sigma$=25).
Setting (1+6) (2+5) (4+3) (5+2) (6+1) (3+4) (ours)
PSNR 32.45 32.56 32.68 32.61 32.51 32.70
SSIM 0.927 0.929 0.931 0.930 0.927 0.931
HAIR for Other Baselines. We have previously posited that HAIR is readily integrable with any image-to-image network architecture. To substantiate this claim, we have implemented HAIR on various established baselines. As depicted in Table Table 6, we selected three efficacious image restoration models—Transweather [1], AirNet [20], and Restormer [12]—for integration with HAIR. Specifically, we have integrated the Classifier at the network’s midpoint for each method and transitioned the subsequent layers to Hypernetworks-based modules. The outcomes shows that our method outperforms PromptIR as a plug-in-and-play module, underscoring the universality and potency of our approach.
Table 6: HAIR for different baselines architectures. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the full RGB images. Best performances are highlighted.
Method Architecture Params Dehazing Deraining Denoising Deblurring Low-Light Average
[…]
Transweather [1] Attention 38M 21.32 .885 29.43 .905 29.00 .841 25.12 .757 21.21 .792 25.22 .836
Transweather+PromptIR 51M 22.89 .920 29.79 .913 29.95 .877 25.74 .781 23.02 .849 26.28 .868
Transweather+HAIR 42M 23.66 .935 32.34 .947 29.96 .875 26.33 .802 23.16 .858 27.09 .884
AirNet [20] CNN 9M 21.04 .884 32.98 .951 30.91 .882 24.35 .781 18.18 .735 25.49 .847
AirNet+PromptIR 14M 21.34 .883 33.52 .953 30.92 .882 24.37 .786 18.18 .737 25.67 .848
AirNet+HAIR 10M 22.15 .899 34.56 .957 30.94 .884 25.44 .792 18.24 .740 26.27 .854
Restormer [12] CNN 26M 24.09 .927 34.81 .962 31.49 .884 27.22 .829 20.41 .806 27.60 .881
Restormer+PromptIR 36M 30.61 .974 36.17 .973 31.25 .887 27.93 .851 22.89 .842 29.77 .905
Restormer+HAIR 29M 30.62 .978 38.11 .981 31.49 .891 28.52 .874 23.12 .847 30.37 .914
5 Conclusion
This paper introduces HAIR, a novel plug-and-play Hypernetworks-based module capable of being easily integrated and adaptively generating parameters for different networks based on the input image. Our method comprises two main components: the Classifier and the Hyper Selecting Net (HSN). Specifically, the Classifier is a simple image classification network with Global Average Pooling, designed to produce a Global Information Vector (GIV) that encapsulates the global information from the input image. The HSN functions as a Fully-connected Neural Network, receiving the GIV and outputting parameters for the corresponding modules. Extensive experiments indicate that HAIR can significantly enhance the performance of various image restoration architectures at a low cost, without necessitating any changes to their logical structures. By incorporating HAIR into the widely recognized Restormer architecture, we have achieved State-Of-The-Art performance on a range of image restoration tasks. The potential for further exploitation of data-conditioned Hypernetworks in image-to-image tasks is substantial, considering their robust adaptability to diverse input images based on their content.