### v4

# Jin Cao1, Yi Cao2, Li Pang1, Deyu Meng1, Xiangyong Cao1

Abstract
Image restoration aims to recover a high-quality clean image from its degraded version. Recent progress in image restoration has demonstrated the effectiveness of All-in-One image restoration models in addressing various unknown degradations simultaneously. However, these existing methods typically utilize the same parameters to tackle images with different types of degradation, forcing the model to balance the performance between different tasks and limiting its performance on each task. To alleviate this issue, we propose HAIR, a Hypernetworks-based All-in-One Image Restoration plug-and-play method that generates parameters based on the input image and thus makes the model to adapt to specific degradation dynamically. Specifically, HAIR consists of two main components, i.e., Degradation-Aware Classifier (DAC) and Hyper Selecting Net (HSN). DAC is a simple image classification network used to generate a Global Information Vector (GIV) that contains the degradation information of the input image, and the HSN is a simple fully-connected neural network that receives the GIV and outputs parameters for the corresponding modules. Extensive experiments demonstrate that HAIR can significantly improve the performance of existing image restoration models in a plug-and-play manner, both in single-task and All-in-One settings. Notably, our proposed model Res-HAIR, which integrates HAIR into the well-known Restormer, can obtain superior performance compared with State-Of-The-Art (SOTA) methods. Moreover, we theoretically demonstrate that to achieve a given small enough error, our proposed HAIR requires fewer parameters in contrast to mainstream embedding-based All-in-One methods.
1 Introduction
Figure 1: Comparisions between our method and previous methods in the inference stage. (a) Previous All-in-One image restoration methods. These methods utilize a single model with fixed parameters to tackle different degradations. (b) Our proposed HAIR. Given a certain degraded image, we use Hypernetworks to generate the dynamic parameters for the decoder, and finally obtain the restored image. Note that "dynamic" and "fixed" in this paper are specially for the main networks, as discussed in Sec. 8.1.
(a) AirNet
(b) PromptIR
(c) HAIR (ours)
(d) HAIR (ours)
Figure 2: Comparison of tSNE plots for the degradation embeddings between previous methods and our HAIR (i.e. the GIVs). Each distinct color represents a unique degradation type. As shown in (c), our HAIR excels not only in recognizing various degradation types, such as noise, rain, and haze, but also in distinguishing between the same type of degradation at varying intensities, e.g. noise with different standard deviations. Even when confronted with composite degradations not encountered during training, HAIR can also accurately discriminate them, i.e. the GIVs for these composite cases located midway between the GIVs of their constituent degradations, as illustrated in (d).
Image restoration is an important task in the field of computer vision, aiming to reconstructing high-quality images from their degraded states. The presence of adverse conditions such as noise, haze, or rain can severely diminish the effectiveness of images for a variety of applications, such as autonomous navigation [48, 8], augmented reality [17, 42, 11]. Therefore, developing robust image restoration methods is of great importance. The use of deep learning in this domain has made remarkable progress, as evidenced by a suite of recent methodologies [64, 29, 6, 45, 25, 67, 60, 7, 28, 56, 13]. Nonetheless, the predominant approach in current research is to employ task-specific models, each tailored to address a particular type of degradation [66, 29, 60, 7, 6]. This tailored approach, while precise, presents a constraint in terms of universality, as it restricts the applicability of models to scenarios with varied or unknown degradations [57, 58, 39]. To overcome this limitation, many researchers have focused on developing All-in-One image restoration models recently. These models are designed to tackle various degradations using one single model. Pioneering efforts in this area [27, 59, 48, 30, 14, 38, 62, 56, 13] have utilized a variety of advanced techniques such as contrastive learning [27], meta-learning [62], visual prompting methods [38, 50, 13, 28, 9]. These approaches have undoubtedly made substantial contributions to this field.
However, these All-in-One models share a common drawback, i.e. they rely on a single model with fixed parameters to address various degradations. This one-size-fits-all method can hinder the model’s effectiveness when dealing with multiple degradations simultaneously. For example, when viewed through the lens of frequency domain analysis, haze is characterized as low-frequency noise, in contrast to rain, which is considered high-frequency interference. An effective dehazing model acts as a low-pass filter, preserving high-frequency details, whereas deraining requires the opposite—enhancing the high-frequency components. Consequently, a model must balance these conflicting demands of different degradations, thus limiting its performance on each task. Further details on this issue are provided in Sec. 7.1.
To mitigate the aforementioned issue, we propose a Hypernetworks-based All-in-One Image Restoration method (HAIR) in this paper. The core idea of HAIR is to generate the weight parameters based on the input image, and thus can dynamically adapt to different degradation information. HAIR employs Hypernetworks [19], a trainable neural network, to take the degradation information from the input image and produce the corresponding parameters. Specifically, for a given unknown degraded image, we first utilize a Degradation-Aware Classifier, similar to those used in image classification networks, to obtain its Global Information Vector, which contains crucial discriminative information about various types of image degradation (as shown in Fig. 2 (c-d)). This vector is then used to generate the necessary parameters, as illustrated in Fig. 1. With these dynamically parameterized modules, we ultimately achieve the restored image. In brief, our contributions include:
  • We propose HAIR, a novel Hypernetworks-based All-in-One image restoration method that is capable of dynamically generating parameters based on the degradation information of input image. HAIR consists of two components, i.e. Degradation Aware Classifier and Hyper Selecting Nets, both of which function as plug-in-and-play modules. Extensive experiments indicate that HAIR can obviously improve the performance of existing models in a plug-and-play manner, both in single-task and All-in-One settings.
  • By incorporating HAIR into Restormer [60], we propose a new All-in-One model, i.e. Res-HAIR. To the best of our knowledge, our method is the first to apply data-adaptive Hypernetworks to All-in-One image restoration models. Extensive experiments validate that the proposed Res-HAIR can achieve superior or comparable performance compared with current state-of-the-art methods across a variety of image restoration tasks.
  • We theoretically prove that, for a given small enough error threshold $\epsilon$ in image restoration tasks, HAIR requires fewer parameters compared to mainstream embedding-based All-in-One methods like [27, 38, 9].
2 Related Works
2.1 All-in-One Image Restoration
While single degradation methods do achieve great success [45, 25, 67, 29, 60, 7], All-in-One image restoration, which aims to utilize a single deep restoration model to tackle multiple types of degradation simultaneously without prior information about the degradation of the input image, has gained more attention recently  [48, 8, 23, 38, 62, 56]. The pioneer work AirNet [27] achieves All-in-One image restoration using contrastive learning to extract degradation representations from corrupted images. IDR [62] decomposes image degradations into their underlying physical principles, achieving All-in-One image restoration through a two-stage process based on meta-learning. With the rise of LLM [68], prompt-based learning has also emerged as a promising direction in image restoration tasks [38, 50, 28]. They typically produce a prompt embedding for each input image based on their content, then inject this prompt into the model to restore the image, which is essentially a conditional embedding. Among them, some works like DA-CLIP [32] and InstructIR [9] insightfully leverage pre-trained large-scale text-vision models to produce the visual prompts. Next, some methods like DaAIR [56] and AMIR [54] utilize routing techniques, which leverage multiple experts within the model to handle different degradation types or tasks by directing the input data along specialized pathways, to achieve adaptive image restoration. However, the output of these multiple experts is also essentially a kind of conditional embedding. Despite the success these methods have achieved, most of them typically use the same parameters for distinct degradations and only inject the degradation information as conditional embeddings into the model, which forces the model to balance different degradations, and thus impair the performance of the models. Tian et al. [47] recently employed low-rank matrix decomposition for weight modulation, but their method relies on pre-trained, task-specific weights and prior degradation information, lacking the dynamic parameter generation capability and adaptibility. In contrast, our proposed HAIR can dynamically generate specific parameters for the given degraded image using a hypernetwork and thus making the model adapt to unknown degradations better.
2.2 Data-conditioned Hypernetworks
Hypernetworks [19] are a class of neural networks designed to generate weights (parameters) for other networks. They can be classified into three types, i.e task-conditioned, data-conditioned, and noise-conditioned hypernetworks [5]. Among these methods, data-conditioned hypernetworks are particularly noteworthy for their ability to generate weights contingent upon the distinctive features of the input data. This capability allows the network to dynamically adapt its behaviour to specific input patterns or characteristics, fostering flexibility and adaptability within the model. Consequently, this results in enhanced generalization and robustness. Data-conditioned hypernetworks have been applied in many computer vision tasks, e.g., semantic segmentation [37] and image editing [2]. Despite previous attempts to integrate hypernetworks into image restoration [1, 14], these have primarily leveraged task-conditioned hypernetworks, which take a task embedding (like an embedding of derain) as the input and output the weights. They have two main drawbacks: (1) They require prior knowledge of the input image’s degradation type. (2) They cannot dynamically generate weights based on image content. Although Klocek et al. [24] tries to use Data-conditional Hypernetworks in Super-Resolution, it direcly maps a coordinate to a pixel, which is unreliable. To the best of our knowledge, our work is the first to introduce data-conditioned hypernetworks into the domain of All-in-One image restoration.
Figure 3: The overall framework of our proposed Res-HAIR. Res-HAIR is built by integrating our HAIR into the popular Restormer [60]. HAIR contains two modules, i.e. Degradation-Aware Classifier (DAC) and Hyper Selecting Net (HSN). The DAC is used to yield a Global Information Vector (GIV) $\mathbf{V_{g}}$ from the high-level feature $\mathbf{F_{g}}$ containing the degradation information of the input image. HSN dynamically generates Transformer Block weights based on GIV, tailoring restoration to each image’s unique degradation.
3 Method
In this section, we first introduce our proposed hypernetwork-based All-in-One image restoration module (HAIR) in detail. Then, we propose our All-in-One method (Res-HAIR) by integrating HAIR into the popular Restormer [60]. Finally, we provide a brief theoretical analysis on the proposed HAIR.
3.1 Proposed HAIR Module
In this subsection, we introduce HAIR, a hypernetworks-based All-in-One image restoration module designed to dynamically generate parameters for an image-to-image network based on the input image’s features in a plug-and-play manner. HAIR comprises two main components: the Degradation-Aware Classifier (DAC) and the Hyper Selecting Net (HSN). Next, we will introduce the two components in detail.
3.1.1 Degradation-Aware Classifier (DAC)
As outlined in Sec. 1, our approach to extracting degradation information involves designing a simple DAC akin to those used in image classification tasks. Taking Fig. 3 as an example, we start by entering the feature $\mathbf{F_{g}}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 8C}$ into the backbone, which contains multiple ResNet blocks [20] followed by downsampling ×$\times 2$. This process is aimed at downsizing the spatial resolution while distilling the essential information, shown as follows:
$\mathbf{F_{g}^{1}}=\text{Downsampling}(\text{ResNetBlock}(\mathbf{F_{g}})),$ (1)
$\mathbf{F_{g}^{t}}=\text{Downsampling}(\text{ResNetBlock}(\mathbf{F_{g}^{t-1}}))),$
where $t=1,2,3$. After three iterations, we obtain $\mathbf{F_{g}^{3}}\in\mathbb{R}^{\frac{H}{64}\times\frac{W}{64}\times 2C}$, which serves as the input of DAC for generating the global information vector (GIV) $\mathbf{V_{g}}\in\mathbb{R}^{2C}$ that captures the degraded information. Specifically,
$\mathbf{V_{g}}=\text{GAP}(\mathbf{F_{g}^{3}}),$ (2)
where GAP denotes Global Average Pooling. The inclusion of GAP ensures that $\mathbf{V_{g}}$ remains a one-dimensional vector of consistent size, irrespective of the input image’s resolution. Unlike traditional image classification networks, there is no requirement for a Softmax layer here, as its application would confine the values of $\mathbf{V_{g}}$ within the range [0,1], potentially restricting the parameter generation process. Moreover, the backbone can be substituted with other image classification networks, e.g., VGG [43] and Inception [44], with little negative impact on performance since degradation-type classification (discrimination) is a relatively simple task.
3.1.2 Hyper Selecting Net (HSN)
After obtaining GIV $\mathbf{V_{g}}$ from the DAC, it is then fed into our data-conditioned hypernetworks, i.e. Hyper Selecting Net (HSN), to generate weights for the corresponding modules. Let’s still take Fig. 3 as an example. Given $\mathbf{V_{g}}\in\mathbb{R}^{2C}$, a Selecting Vector $\mathbf{V_{s}}\in\mathbb{R}^{N}$ is initially computed as:
$\mathbf{V_{s}}=\sigma(\mathbf{FCNN}(\mathbf{V_{g}})),$ (3)
where $\sigma$ is a Softmax operation and $\mathbf{FCNN}$ denotes a simple fully-connected neural network. Unlike $\mathbf{V_{g}}$, $\mathbf{V_{s}}$ will be used to directly generate the parameters, so we need a $\sigma$ to make sure $\mathbf{V_{s}}$ is positive and $\in[0,1]$, thus making parameter generation more stable. Subsequently, the parameters
$\mathbf{w}=\sum_{i=1}^{N}\mathbf{V_{s}}^{i}\mathbf{W}_{i}.$ (4)
In this formula, $\mathbf{V_{s}}^{i}$ represents the $i$-th element of $\mathbf{V_{s}}$. The matrix $\mathbf{W}\in\mathbb{R}^{N\times P}$, referred to as the Weight Box, comprises $\mathbf{W}_{i}\in\mathbb{R}^{P}$ as its $i$-th row. The hyperparameter $N$ influences the total number of parameters, with $P$ being the count of parameters required for one corresponding module (i.e. one Transformer Block in this example). With $\mathbf{w}$ determined, it is then used as the parameters of the corresponding module. So the operation of the "HyperTrans Block" in Fig. 3 can be represented as:
$\mathbf{x}^{\prime}=\mathbf{Transformer\_Block}(\mathbf{x};\mathbf{w}).$ (5)
Here, $\mathbf{x}$ is the input and $\mathbf{x}^{\prime}$ is the output. As the Transformer Block (see Sec. 9) is based on convolution, $\mathbf{w}\in\mathbb{R}^{P}$ is reshaped into four-dimensional tensors to serve as convolution kernels for the Transformer Block. To reduce the number of parameters, the transformer blocks at the same decoder level share one weight box, and each transformer block is independently equipped with its own FCNN.
3.2 Proposed Res-HAIR Model
Integrating HAIR into an existing image-to-image network involves a simple two-step process: (1) Insert the DAC at the junction between the first and second halves of the network to produce a Global Information Vector (GIV) from the features of the first half. (2) Incorporate HSNs into the second half of the network, using the GIV to generate weights for all modules within this section dynamically. In this work, we integrate our proposed HAIR module into the popular image restoration model Restormer [60], and propose our All-in-One model Res-HAIR. Although we only use the example of integrating HAIR with Restormer to illustrate the process, this example is representative, and integration with other networks follows in a virtually identical way.
Overall Architecture. As shown in Fig. 3, our network architecture is consistent with Restormer [60]. Given a certain degraded image input $\mathbf{X}\in\mathbb{R}^{H\times W\times 3}$, Res-HAIR utilizes a $3\times 3$ convolution to transform $\mathbf{X}$ into feature embeddings $\mathbf{F_{0}}\in\mathbb{R}^{H\times W\times C}$, where $C$ denotes the number of channels. These feature embeddings are then proceeded through a 4-level hierarchical encoder-decoder, resulting in deep features $\mathbf{F_{r}}$. Each level of the encoder-decoder incorporates several Transformer blocks, with an increasing number from the top to the bottom level, ensuring computational efficiency. The left three blue "Transformer Blocks" function as an Encoder, designed to extract features from $\mathbf{F_{0}}$ and ultimately produce a global feature map $\mathbf{F_{g}}\in\mathbb{R}^{\frac{W}{8}\times\frac{H}{8}\times 8C}$ with a large receptive field. $\mathbf{F_{g}}$ is then input into the DAC to yield a Global Information Vector (GIV) $\mathbf{V_{g}}\in\mathbb{R}^{2C}$. The right four orange "HyperTrans Blocks" operate as a Decoder, aiming to adaptively fuse features at each Decoder level based on degradation, culminating in a restored image $\hat{\mathbf{X}}$. Specifically, at each Decoder level, the HyperTrans Block receives two inputs, i.e. the input feature and the GIV $\mathbf{V_{g}}$. $\mathbf{V_{g}}$ is fed into the Hyper Selecting Nets to generate weights for the corresponding Transformer Blocks, which are then applied to the input feature to produce the output. It is important to note that all HyperTrans Blocks use the same $\mathbf{V_{g}}$ derived from $\mathbf{F_{g}}$. Since the weights in the Decoder are generated based on $\mathbf{V_{g}}$, which contains the degradation information of $\mathbf{X}$, our method can thus adapt to different degraded images.
3.3 Theoretical Analysis
In this subsection, we first interpret the generated weights from a selection perspective and then provide a brief theoretical analysis of model complexity.
3.3.1 Hypernetworks from a Selection Perspective
Considering the Weight Box as part of the FCNN, the direct generation of weights may still appear somewhat opaque. To clarify this process, we first introduce a proposition before diving into the detailed explanation.
Proposition 1.
(Convolution operations exhibit the distributive law over addition) Let $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$ be the input feature, and let $w_{i},\ i=1,2,\cdots,n$ represent the convolution kernels. The law is mathematically expressed as:
$\mathbf{x}*(\sum_{i=1}^{n}w_{i})=\sum_{i=1}^{n}(\mathbf{x}*w_{i})$ (6)
where ’*’ denotes the standard 2-dimensional convolution.
The following formula then becomes evident:
$\mathbf{x}*\mathbf{w}=\mathbf{x}*\left(\sum_{i=1}^{N}\mathbf{V_{s}}^{i}\mathbf{W}_{i}\right)=$ (7)
$\sum_{i=1}^{N}\left(\mathbf{x}*(\mathbf{V_{s}}^{i}\mathbf{W}_{i})\right)=\sum_{i=1}^{N}\mathbf{V_{s}}^{i}\left(\mathbf{x}*\mathbf{W}_{i}\right),$
where $\sum_{i}^{N}\mathbf{V_{s}}^{i}=1$ holds due to the property of Softmax operation. Essentially, the convolution with the generated weights is equivalent to a weighted sum of convolutions between the input $\mathbf{x}$ and each kernel $\mathbf{W}_{i}$ in the Weight Box. This process can be seen as a process of selecting convolution kernels. If we view each $\mathbf{W}_{i}$ as an expert, it’s also like the Mixture-of-Experts pattern used in [56]. For instance, with a Weight Box $\mathbf{W}\in\mathbb{R}^{2\times P}$ that includes a low-pass $\mathbf{W}_{1}$ and a high-pass $\mathbf{W}_{2}$, the rainy input $\mathbf{x}_{1}$ would ideally have a larger $\mathbf{V_{s}}^{1}$ to filter out high-frequency noise, while a smaller $\mathbf{V_{s}}^{2}$ would help retain details. In contrast, for a hazy input $\mathbf{x}_{2}$, a larger $\mathbf{V_{s}}^{2}$ would be necessary to mitigate low-frequency haze. Although real-world degradations can be more complex, our HSN can adaptively select the appropriate weights. This insight into the selection process helps us understand HAIR, i.e., HSN produces a tailored Selecting Vector and adaptively chooses the most suitable convolution kernel for each input. Given that core operations such as convolution and matrix multiplication follow the distributive law over addition, HAIR is universal and can be integrated into architectures like Transformer [49] and Mamba [18].
3.3.2 A Brief Analysis of Model Complexity
In Sec. 2.2, we claim that our Hypernetworks-based method can be more efficient than conditional embedding-based methods such as AirNet [27] and PromptIR [38]. This section will provide a simple proof for this point. In the context of All-in-One image restoration, we aim to find a function $f:\mathbb{R}^{3HW}\xrightarrow{}\mathbb{R}^{3HW}$ to map a degraded image $\mathbf{X}$ to a restored image $\hat{\mathbf{X}}$ by minimizing the distance between $\hat{\mathbf{X}}$ and the ground truth $\mathbf{Y}=y(\mathbf{X})$ [^1] For the sake of discussion, the following tensors are generally regarded as flattened one-dimensional vectors.. Mainstream methods typically utilize a network $g(\mathbf{X},e(\mathbf{X}))$ to learn the mapping, where $e(\mathbf{X})\in\mathbb{R}^{k}$ contains the degradation embedding of the input image, such as the prompt in PromptIR [38]. These methods send $e(\mathbf{X})$ together with $\mathbf{X}$ into the network to get $\hat{\mathbf{X}}$. For our HAIR, the network can be formulated as $h(\mathbf{X};\theta(e(\mathbf{X})))$, where $\theta$ is a function that maps ${e(\mathbf{X})}$ into the parameters of $h$. We define the distance between functions as
$d(g,y)=\min_{g}\max_{\mathbf{X}}\|g(\mathbf{X},e(\mathbf{X}))-y(\mathbf{X})\|_{\infty}.$ (8)
Then we have the following 2 theorems.
Theorem 1.
For conditional embedding-based All-in-One image restoration methods $g(\mathbf{X},e(\mathbf{X}))$ like PromptIR [38], to achieve error $d(g,y)\leq\epsilon$, the complexity (number of parameters) of the corresponding model is at least $N_{g}=\Omega\left(\epsilon^{-\min(3HW+k,6HW)}\right)$, where $k$ is the dimensionality of the embedding vector. Typically $k=\Omega(HW)$, $k\geq 3HW$.
Theorem 2.
For Hypernetworks based methods $h(\mathbf{X};\theta(e(\mathbf{X})))$ like our HAIR, to achieve error $d(h,y)\leq\epsilon$, the complexity of the corresponding model is $N_{\theta}=O(\epsilon^{-3HW/r})$, where $r\geq 1$.
The two theorems demonstrate that to achieve the same error $\epsilon$, our proposed HAIR requires fewer parameters than conditional embedding-based methods as long as $\epsilon$ is small enough. Please refer to [15] and Sec. 6 in supplementary materials for detailed proof of Theorem 1 and 2.
4 Experiments
The experimental settings follow previous research on general image restoration [62, 38] in two different configurations: (a) All-in-One and (b) Single task. In the All-in-One configuration, a singular model is trained to handle multiple types of degradation, encompassing both three and five unique degradations. In contrast, the Single-task configuration involves training individual models dedicated to specific restoration tasks.
4.1 Experimenting Preparation
Implementation Details. Our Res-HAIR method is designed to be end-to-end trainable, eliminating the need for pre-training any of its components. The architecture comprises a 4-level encoder-decoder structure, with each level equipped with a varying number of Transformer blocks. Specifically, the number of blocks increases progressively from level-1 to level-4, following the sequence [4, 6, 6, 8]. HyperTrans Blocks are employed throughout level-4 and the decoding stages of levels 1-3. Additionally, the Weight Box parameter $N$ is set according to the sequence [5, 7, 7, 9] for each respective level. In the All-in-One setting, the model is trained with a batch size of 32, while in the single-task setting, a batch size of 8 is used. The network optimization is guided by an L1 loss function, employing the AdamW optimizer [31] with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate is set to $42e-$ 4 for the initial 150 epochs and then changed to $52e-$ 5 for the final 10 epochs. To enhance the training data, input patches of size 128 $\times$ 128 are utilized, with random horizontal and vertical flips applied to the images to augment the dataset.
Datasets. We follow previous approaches [27, 38, 62] for our All-in-One and single-task experiments, using these benchmark datasets. For single-task image denoising, we use the BSD400 [3] and WED [33] datasets, adding Gaussian noise with levels $\sigma\in\{15,25,50\}$ to generate training images. Testing is performed on the BSD68 [34] and Urban100 [21] datasets. For deraining, we employ the Rain100L dataset [53]. Dehazing experiments utilize the SOTS dataset [26]. Deblurring and low-light enhancement tasks use the GoPro [36] and LOL-v1 [51] datasets, respectively. To create a comprehensive model for all tasks, we combine these datasets and train them in a setting that covers three or five types of degradations. For single-task models, training is conducted on the respective dataset.
Table 1: Comparison to state-of-the-art on three degradations. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the full RGB images with $(^{\ast})$ denoting methods that are not blind (need prior information of the degradation). Best and second best performances are highlighted.
Method Params. Dehazing Deraining Denoising Average
[…]
BRDNet [46] - 23.23 .895 27.42 .895 32.26 .898 29.76 .836 26.34 .693 27.80 .843
LPNet [16] - 20.84 .828 24.88 .784 26.47 .778 24.77 .748 21.26 .552 23.64 .738
FDGAN [12] - 24.71 .929 29.89 .933 30.25 .910 28.81 .868 26.43 .776 28.02 .883
DL [14] 2M 26.92 .931 32.62 .931 33.05 .914 30.41 .861 26.90 .740 29.98 .876
MPRNet [59] 16M 25.28 .955 33.57 .954 33.54 .927 30.89 .880 27.56 .779 30.17 .899
Restormer [60] 26M 29.92 .970 35.56 .969 33.86 .933 31.20 .888 27.90 .794 31.68 .910
AirNet [27] 9M 27.94 .962 34.90 .967 33.92 .933 31.26 .888 28.00 .797 31.20 .910
PromptIR [38] 36M 30.58 .959 36.37 .972 33.98 .933 31.31 .888 28.06 .799 32.06 .913
InstructIR [9] 16M 30.22 .959 37.98 .978 34.15 .933 31.52 .890 28.30 .804 32.43 .913
Res-HAIR (ours) 29M 30.98 .979 38.59 .983 34.16 .935 31.51 .892 28.24 .803 32.70 .919
4.2 Results
All-in-One: (1) Three degradations. We conducted a comparative evaluation of our All-in-One image restoration model against several state-of-the-art specialized methods, including BRDNet [46], LPNet [16], FDGAN [12], DL [14], MPRNet [59], AirNet [27], PromptIR [38], InstructIR [9] and DaAIR [56]. Each of these methods was trained to handle the tasks of dehazing, deraining, and denoising. As illustrated in Tab. 1, our approach can significantly outperform other competing methods, e.g. an average improvement of $0.64$ dB over PromptIR. Our method notably outperforms existing benchmarks on SOTS and Rain100L datasets, by exceeding the performance of the previous best methods by margins of $0.4$ dB and $0.61$ dB respectively. (2) Five Degradations. Following recent studies [62, 9], we have extended our approach to deblurring and low-light image enhancement, thus extending the three-degradation setting to a more complex five-degradation setting. As demonstrated in Tab. 2, our method excels by learning specialized models for each degradation type while effectively capturing the shared characteristics between tasks. It achieves an average improvement of $2.03$ dB over PromptIR and $0.82$ dB over the non-blind method InstructIR, establishing a new benchmark for state-of-the-art performance across all five benchmarks. Notably, our method significantly outperforms our baseline (i.e. Restormer) in various tasks. This comparison highlights the strong capability of our method as a versatile plug-in-and-play module, enhancing the performance of the existing model with a small amount of integration complexity, e.g. only adding 3M parameters compared with Restormer.
Table 2: Comparison to state-of-the-art on five degradations. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the full RGB images with $(^{\ast})$ denoting methods that are not blind (need prior information of the degradation type). Best and second best performances are highlighted.
Method Params. Dehazing Deraining Denoising Deblurring Low-Light Average
[…]
NAFNet [6] 17M 25.23 .939 35.56 .967 31.02 .883 26.53 .808 20.49 .809 27.76 .881
DGUNet [35] 17M 24.78 .940 36.62 .971 31.10 .883 27.25 .837 21.87 .823 28.32 .891
SwinIR [29] 1M 21.50 .891 30.78 .923 30.59 .868 24.52 .773 17.81 .723 25.04 .835
Restormer [60] 26M 24.09 .927 34.81 .962 31.49 .884 27.22 .829 20.41 .806 27.60 .881
DL [14] 2M 20.54 .826 21.96 .762 23.09 .745 19.86 .672 19.83 .712 21.05 .743
Transweather [48] 38M 21.32 .885 29.43 .905 29.00 .841 25.12 .757 21.21 .792 25.22 .836
TAPE [30] 1M 22.16 .861 29.67 .904 30.18 .855 24.47 .763 18.97 .621 25.09 .801
AirNet [27] 9M 21.04 .884 32.98 .951 30.91 .882 24.35 .781 18.18 .735 25.49 .847
IDR [62] 15M 25.24 .943 35.63 .965 31.60 .887 27.87 .846 21.34 .826 28.34 .893
PromptIR [38] 36M 30.61 .974 36.17 .973 31.25 .887 27.93 .851 22.89 .842 29.77 .905
InstructIR [9] 16M 27.00 .951 36.80 .973 31.39 .888 29.73 .890 22.83 .836 29.55 .908
Res-HAIR (ours) 29M 30.62 .978 38.11 .981 31.49 .891 28.52 .874 23.12 .847 30.37 .914
Single-Degradation Results. To evaluate the effectiveness of our proposed method, we provide results in Tab. 3, which show the performance of individual instances of our method trained under a single degradation protocol. Specifically, the single-task variant dedicated to deraining consistently achieves higher performance than PromptIR and InstructIR by margins of $1.96$ dB and $1.02$ dB, respectively. When applied to image denoising, our method also demonstrates superiority over the aforementioned approaches, with an average improvement of $0.21$ dB and $0.47$ dB, respectively. These results underscore the significant performance improvements delivered by our method.
Table 3: Comparison to state-of-the-art for single degradations. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the full RGB images. Best and second best performances are highlighted. Denoising results are from Urban100  [21].
(a) Dehazing
Method SOTS
DehazeNet  [4] 22.46 .851
MSCNN  [41] 22.06 .908
EPDN  [40] 22.57 .863
FDGAN [12] 23.15 .921
Restormer [60] 30.87 .969
AirNet [27] 23.18 .900
PromptIR [38] 31.31 .973
InstructIR [9] 30.22 .959
Res-HAIR (ours) 31.68 .980
(b) Deraining
Method Rain100L
DIDMDN [61] 23.79 .773
UMR [55] 32.39 .921
SIRR [52] 32.37 .926
MSPFN [22] 33.50 .948
Restormer [60] 36.74 .978
AirNet [27] 34.90 .977
PromptIR [38] 37.04 .979
InstructIR [9] 37.98 .978
Res-HAIR (ours) 39.00 .985
(c) Denoising
Method $\sigma$=15 $\sigma$=25 $\sigma$=50
CBM3D [10] 33.93 .941 31.36 .909 27.93 .833
DnCNN [63] 32.98 .931 30.81 .902 27.59 .833
IRCNN [64] 27.59 .833 31.20 .909 27.70 .840
FFDNet  [65] 33.83 .942 31.40 .912 28.05 .848
BRDNet [46] 34.42 .946 31.99 .919 28.56 .858
AirNet [27] 34.40 .949 32.10 .924 28.88 .871
PromptIR [38] 34.77 .952 32.49 .929 29.39 .881
InstructIR [9] 34.12 .945 31.80 .917 28.63 .861
Res-HAIR (ours) 34.93 .953 32.70 .931 29.65 .885
Visual Results. The visual results captured under three degradation scenarios are presented in Fig. 4. These results clearly demonstrate the superiority of our method in terms of visual quality. In the denoising task at $\sigma=25$, Res-HAIR retains more fine details compared to other methods; in the deraining task, Res-HAIR effectively removes all rain streaks, whereas the compared methods contain obvious rain streaks; in the dehazing task, Res-HAIR may not closely resemble the ground-truth, yet it provides the most visually pleasing outcome and even clears the original haze present in the ground-truth. Overall, these visual observations confirm the efficacy of our approach in enhancing image quality across different degradation conditions.
Figure 4: Visual comparison on various degradation settings. Zoom for better visual effects.
4.3 Ablation Studies
We have done several ablation studies to evaluate the effectiveness of our proposed HAIR.
Effectiveness of DAC and HSN. As detailed in Tab. 5, we study the impact of the proposed modules in four settings: (a) The model aligns with the Restormer architecture [60]. (b) In this configuration, the Global Information Vector (GIV) is directly used as conditional embeddings for the Transformer Blocks in the Decoder, rather than for weight generation. (c) The GIVs are designated as independently trainable parameters, each randomly initialized, with the Decoder’s Transformer Blocks having their distinct GIVs. (d) This setup incorporates both components. The results demonstrate the indispensability of both components. With the addition of only 3M parameters and no change to the logical structure, Res-HAIR outperforms Restormer by 1.7 dB in PSNR, demonstrating its simplicity and effectiveness.
Table 4: HAIR for different baseline architectures. PSNR (dB, $\uparrow$) and SSIM ($\uparrow$) metrics are reported on the full RGB images. Best performances are highlighted.
Method Main Operation Params Dehazing Deraining Denoising Deblurring Low-Light Average
[…]
Transweather [48] Self-Attention 38M 21.32 .885 29.43 .905 29.00 .841 25.12 .757 21.21 .792 25.22 .836
Transweather [48]+PromptIR [38] 51M 22.89 .920 29.79 .913 29.95 .877 25.74 .781 23.02 .849 26.28 .868
Transweather [48]+HAIR 42M 23.66 .935 32.34 .947 29.96 .875 26.33 .802 23.16 .858 27.09 .884
AirNet [27] Convolution 9M 21.04 .884 32.98 .951 30.91 .882 24.35 .781 18.18 .735 25.49 .847
AirNet [27]+PromptIR [38] 14M 21.34 .883 33.52 .953 30.92 .882 24.37 .786 18.18 .737 25.67 .848
AirNet [27]+HAIR 10M 22.15 .899 34.56 .957 30.94 .884 25.44 .792 18.24 .740 26.27 .854
Restormer [60] Convolution 26M 24.09 .927 34.81 .962 31.49 .884 27.22 .829 20.41 .806 27.60 .881
Restormer [60]+PromptIR [38] 36M 30.61 .974 36.17 .973 31.25 .887 27.93 .851 22.89 .842 29.77 .905
Restormer [60]+HAIR 29M 30.62 .978 38.11 .981 31.49 .891 28.52 .874 23.12 .847 30.37 .914
Position of DAC and HyperTrans Blocks. In our approach, the DAC is positioned after the first three Encoder levels, with the subsequent four Decoder levels employing HyperTrans Blocks, constituting a 3+4 configuration. The rationale for such design is shown in Tab. 6. As can be seen, the optimal placement for the DAC is at the network’s midpoint. This positioning takes advantage of the expansive receptive field of the features post the network’s halfway point, while allowing the HSN to dynamically generate parameters for the remaining modules. Furthermore, it is crucial to strike a balance between the receptive field size of the features fed into the DAC and the count of HyperTrans Blocks utilized, ensuring the model’s efficiency and adaptability.
Table 5: Impact of the existence of DAC and HSN. Results are from single deraining task on Rain100L.
Setting Prams. DAC HSN PSNR SSIM
(a) (baseline) 26M 36.74 .978
(b) (with DAC) 27M 36.88 .979
(c) (with HSN) 28M 36.76 .979
(d) (ours) 29M 39.00 .985
Table 6: Impact of position of DAC and HyperTrans Blocks. Results are from Denoise task on Urban100 ($\sigma$=25).
Setting (1+6) (2+5) (4+3) (5+2) (6+1) (3+4) (ours)
PSNR 32.45 32.56 32.68 32.61 32.51 32.70
SSIM 0.927 0.929 0.931 0.930 0.927 0.931
HAIR for Different Baselines. We have previously posited that HAIR is essentially a plug-in-and-play module which is readily integrable with any existing network architecture. To substantiate this claim, we have implemented HAIR on various baselines. As depicted in Tab. 4, we selected three efficacious image restoration models, i.e. Transweather [48], AirNet [27], and Restormer [60] for integration with HAIR. Specifically, we have integrated the DAC at the network’s midpoint for each method and transitioned the subsequent layers to Hypernetworks-based modules. The results show that our HAIR can significantly improve the performance of these baselines. Additionally, our HAIR outperforms PromptIR as a plug-in-and-play module, demonstrating its better application value.
5 Conclusion
This paper introduces HAIR, a novel plug-and-play Hypernetworks-based module capable of being easily integrated and adaptively generating parameters for different networks based on the input image. Our method comprises two main components: the Degradation-Aware Classifier (DAC) and the Hyper Selecting Net (HSN). Specifically, the DAC is a simple image classification network with Global Average Pooling, designed to produce a Global Information Vector (GIV) that encapsulates the global information from the input image. The HSN functions as a fully-connected neural network, receiving the GIV and outputting parameters for the corresponding modules. Extensive experiments indicate that HAIR can significantly enhance the performance of various image restoration architectures at a low cost without necessitating any changes to their logical structures. By incorporating HAIR into the widely recognized Restormer architecture, we have achieved SOTA performance on a range of image restoration tasks.
Supplementary Material
6 Detailed Proof of Theorem 1, Theorem 2
Notations We consider $\mathcal{X}=[-1,1]^{m_{1}}$ and $\mathcal{I}=[-1,1]^{m_{2}}$ and denote, $m:=m_{1}+m_{2}$. Here $\mathcal{X}$ stands for the set of input images $\mathbf{X}$, meanwhile $\mathcal{I}$ refers to the set of degradation information $e(\mathbf{X})$. For a closed set $X\subset\mathbb{R}^{n}$, we denote by $C^{r}(X)$ the linear space of all $r$-continuously differentiable functions $h:X\to\mathbb{R}$ on $X$ equipped with the supremum norm $\|h\|_{\infty}:=\max_{x\in X}\|h(x)\|_{1}$.
Given that vector functions can be complex, we define a scalar function $f^{i}:\mathbb{R}^{M}\xrightarrow{}\mathbb{R}(i=1,\cdots M)$ of a vector function $f:\mathbb{R}^{M}\xrightarrow{}\mathbb{R}^{M}$. Since $f^{i}$ is the $i$-th element of $f$, then we have
$d(g,y)$ $=\min_{g}\max_{\mathbf{X}}\max_{i}|g^{i}(\mathbf{X})-y^{i}(\mathbf{X})|$ (9)
$=\min_{g}\max_{i}\max_{\mathbf{X}}|g^{i}(\mathbf{X})-y^{i}(\mathbf{X})|$
$=\min_{g}\max_{\mathbf{X}}|g^{t(g)}(\mathbf{X})-y^{t(g)}(\mathbf{X})|.$
Since the value range of $i$ is limited, given the function $g$, we can always find an integer $t(g)$ to replace $\max_{i}$. To complete the proof, we need some assumptions.
Assumption 1.
The target function $y^{i}\in\mathcal{W}_{r,3HW},i=1,2,\cdots 3HW$. The Sobolev space $\mathcal{W}_{r,3HW}$ is the set of functions that are $r$-times differentiable ($r\geq 1$) with all $r$-th order derivatives being continuous and bounded by the Sobolev norm $\leq 1$, defined on $\mathbb{R}^{3HW}$.
Intuitively, since $y^{i}$ maps the degraded image to a single pixel of the clean image, a slight disturbance in the degraded image should only bring a slight difference to the single pixel, so $y^{i}$ can be smooth and differentiable, or at least continue. Moreover, the domain and range of $y$ are restricted to $[0,1]^{3HW}$, with most pixels far less than 1, the Sobolev norm of $y_{i}$ can generally be less than 1 and thus this assumption generally holds.
Assumption 2.
For various functions (e.g. network) $g$, the integer $t(g)$ in Eq. 9 remains the same.
In our context, $g$ represents the same network with different weights during training. No matter how the parameters update, the most "difficult" degraded image generally remains the same one, e.g., the one with very severe degradation. Within this "difficult" image, the most "difficult" pixel should also remain the same, e.g. the pixel that varies dramatically from its clean version. In this way, we simplify the vector function $g$ to scalar function $g^{t(g)}$. With the two assumptions and all the assumptions in [15], we can directly use Theorem 3,4 and 5 to analyze the complexity.
Theorem 3.
(The Theorem 2 in [15]) Let $\sigma$ be a universal, piece-wise $C^{1}(\mathbb{R})$ activation function with $\sigma^{\prime}\in BV(\mathbb{R})$ and $\sigma(0)=0$. Let $\mathcal{E}_{\mathbf{e},\mathbf{q}}$ be a neural embedding method. Assume that $\mathbf{e}$ is a class of continuously differentiable neural network $e$ with zero biases, output dimension $k=\mathcal{O}(1)$ and $\mathcal{C}(e)\leq\ell_{1}$ and $\mathbf{q}$ is a class of neural networks $q$ with $\sigma$ activations and $\mathcal{C}(q)\leq\ell_{2}$. Let $\mathbb{Y}:=\mathcal{W}_{1,m}$. Assume that any non-constant $y\in\mathbb{Y}$ cannot be represented as a neural network with $\sigma$ activations. If the embedding method achieves error $d(\mathcal{E}_{\mathbf{e},\mathbf{q}},\mathbb{Y})\leq\epsilon$, then, the complexity of $\mathbf{q}$ is: $N_{\mathbf{q}}=\Omega\left(\epsilon^{-(m_{1}+m_{2})}\right)$.
The notation $BV(\mathbb{R})$ stands for the set of functions of bounded variation,
$BV(\mathbb{R}):=\left\{f\in L^{1}(\mathbb{R})\mid\|f\|_{BV}<\infty\right\}\textnormal{ where, }$, (10)
$\|f\|_{BV}:=\sup\limits_{\begin{subarray}{c}\phi\in C^{1}_{c}(\mathbb{R})\\ \|\phi\|_{\infty}\leq 1\end{subarray}}\int_{\mathbb{R}}f(x)\cdot\phi(x)\;\textnormal{d}x$
Note that a distinct neural network $e$ is not mandatory. For example, the "prompts" in PromptIR [38] are a set of trainable parameters that do not require a separate network to generate them. Yet, the conclusion remains the same even if network $e$ is non-existent.
Theorem 4.
(The Theorem 3 in [15]) In the setting of Theorem 3, except $k$ is not necessarily $\mathcal{O}(1)$. Assume that the first layer of any $q\in\mathbf{q}$ is bounded $\|W^{1}\|_{1}\leq c$, for some constant $c>0$. If the embedding method achieves error $d(\mathcal{E}_{\mathbf{e},\mathbf{q}},\mathbb{Y})\leq\epsilon$, then, the complexity of $\mathbf{q}$ is: $N_{\mathbf{q}}=\Omega\left(\epsilon^{-\min(m,2m_{1})}\right)$.
Theorem 5.
(The Theorem 4 in [15]) Let $\sigma$ be as in Theorem 3. Let $y\in\mathbb{Y}=\mathcal{W}_{r,m}$ be a function, such that, $y_{I}$ cannot be represented as a neural network with $\sigma$ activations for all $I\in\mathcal{I}$. Then, there is a class, $\mathbf{g}$, of neural networks with $\sigma$ activations and a network $f(I;\theta_{f})$ with ReLU activations, such that, $h(x,I)=g(x;f(I;\theta_{f}))$ achieves error $\leq\epsilon$ in approximating $y$ and $N_{\mathbf{g}}=\mathcal{O}\left(\epsilon^{-m_{1}/r}\right)$.
In the realm of image restoration, $m_{1}$ equals $3HW$, and $m_{2}$ equals $k$, where $k$ denotes the dimensionality of the flattened degradation embedding. In our method, $k$ is consistent with the shape of the Global Information Vector (GIV), specifically $2C$, and thus is $\mathcal{O}(1)$. Conversely, in PromptIR [38], $k$ is dynamic and contingent on the resolution of the input image, precluding it from being $\mathcal{O}(1)$. Theorem 5 indicates that the complexity for a Hypernetworks-based method to attain an error of $\epsilon$ is $\mathcal{O}\left(\epsilon^{-3HW/r}\right)$. Theorems 3 and 4 collectively suggest that the complexity for embedding-based methods is at least $\Omega\left(\epsilon^{-\min(3HW+k,6HW)}\right)$. This comparison shows that Hypernetworks-based methods like HAIR may require fewer parameters to reach a given error threshold compared to their embedding-based counterparts.
7 Discussion
7.1 HAIR for Conflicting Degradations.
As previously discussed in the Introduction (Sec. 1), the performance of conventional All-in-One image restoration methods, which rely on a single model with static parameters, can be significantly compromised when dealing with conflicting image degradations. To demonstrate this, we trained models on various combinations of datasets, each representing different types of degradations, and subsequently evaluated these models on their respective benchmarks. The outcomes are presented in Tables 7 and 8.
From a frequency domain analysis perspective, haze is identified as low-frequency noise, whereas rain and Gaussian noise are categorized as high-frequency disturbances. Conflicts arise when the degradations in a combined scenario include both low- and high-frequency noise components. According to Table 7, it is evident that the presence of conflicting degradations, such as the combination of noise and haze or rain and haze, can severely degrade the model’s performance. In contrast, when the combined degradations do not conflict, like the combination of noise and rain, the performance loss is minimal and may even enhance the overall performance.
Table 8 exhibits a similar trend, but with a notably reduced impact from conflicting degradations. This reduction in performance impairment underscores the effectiveness of our proposed Hypernetworks-based approach, which dynamically generates parameters based on the input image’s content. This adaptability allows our method to mitigate the performance loss typically associated with conflicting degradations.
Table 7: Performance of the PromptIR [38], when trained on different combinations of degradation types (tasks) i.e., removal of gaussian noise, rain and haze. Note that the "combination" here stands for combination of Datasets with single degradations instead of composite degradations. "Conflicting" here shows if the combined degradations are conflicting to each other. PSNR/SSIM are reported.
Degradation Conflicting Denoising on BSD68 dataset Deraining on Dehazing on
Noise Rain Haze $\sigma=15$ $\sigma=25$ $\sigma=50$ Rain100 SOTS
- 34.34/0.938 31.71/0.898 28.49/0.813 - -
- - - - 37.04/0.979 -
- - - - - 31.31/0.973
34.26/0.937 31.61/0.895 28.37/0.810 39.32/0.986 -
33.69/0.928 31.03/0.880 27.$7$ 30.09/0.975
- - - 35.12/0.969 30.36/0.973
33.98/0.933 31.31/0.888 28.06/0.799 36.37/0.972 30.58/0.974
Table 8: Performance of the our proposed Res-HAIR, when trained on different combinations of degradation types (tasks) i.e., removal of gaussian noise, rain and haze. PSNR/SSIM are reported.
Degradation Conflicting Denoising on BSD68 dataset Deraining on Dehazing on
Noise Rain Haze $\sigma=15$ $\sigma=25$ $\sigma=50$ Rain100 SOTS
- 34.36/0.938 31.72/0.898 28.50/0.813 - -
- - - - 39.00/0.985 -
- - - - - 31.68/0.980
34.33/0.937 31.66/0.896 28.44/0.811 41.55/0.989 -
34.13/0.935 31.49/0.892 28.22/0.803 - 31.18/0.979
- - - 38.44/0.983 31.22/0.979
34.16/0.935 31.51/0.892 28.24/0.803 38.59/0.983 30.98/0.979
7.2 HAIR for Unseen Composite Degradation.
Since HAIR can accurately discriminate unseen composite degradations, as illustrated in Fig. 2(d), we test Res-HAIR with these settings, as shown in Table 9, 11, 10. The results somehow shows the generalization ability of our proposed method. However, we consider the restoration outcomes for such degradations are not satisfactory enough. The interesting fact is, sometimes HAIR generates GIVs that are intermediate when confronted with composite degradations. This tendency results in weights that are a midpoint between those associated with each individual degradation type. For example, an image with both noise and haze degradations results in weights that are intermediate between the weights for images affected by noise alone and those affected by haze alone, which fails to fully eliminate either degradation and result in a somehow unsatisfactory performance. This insight reveals HAIR’s operational mechanism, highlighting its strategy for weights generation.
Table 9: Results of unseen composite degradation (haze + noise) on SOTS.
Method PSNR SSIM
PromptIR 16.211 0.785
Res-HAIR 16.798 0.802
Table 10: Results of unseen composite degradation (rain + noise) on Rain100L.
Method PSNR SSIM
PromptIR 24.348 0.726
Res-HAIR 24.365 0.729
Table 11: Results of unseen composite degradation (rain + haze) on SOTS.
Method PSNR SSIM
PromptIR 23.784 0.686
Res-HAIR 23.823 0.692
8 Possible Confusions
8.1 Does HAIR Also Utilize Fixed Parameters as Previous Methods Do?
In the Introduction (Sec. 1), we highlighted that our proposed HAIR method differs from previous approaches, which typically employ a static set of parameters to handle various image degradations. In contrast, HAIR employs a data-conditioned Hypernetwork to dynamically generate parameters based on the input image’s content. However, a potential point of confusion arises from the fact that the Hypernetworks in HAIR (i.e., the Hyper Selecting Nets) are indeed fixed during inference, suggesting that HAIR also relies on a set of fixed parameters to address different degradations.
To clarify this confusion, it is crucial to understand our motivation for using Hypernetworks. Our goal is to mitigate the performance loss caused by conflicting degradations. For example, by generating distinct parameters, we aim to enable the model to function either as a low-frequency or high-frequency filter, depending on the input image’s requirements. As illustrated in Fig. 1, the primary focus should be on the main networks that directly process the image.
Therefore, when we refer to "fixed parameters" and "dynamic parameters," we are referring to the parameters of the main network that interacts with the image, not the Hypernetworks responsible for parameter generation, which do not directly engage with the input image. The point lies in the dynamic adaptation of the main network’s parameters to the specific characteristics of the input image, which is the innovative aspect of HAIR.
9 More Implementation Details
Figure 5: Overview of the Transformer block used in the Res-HAIR framework. The Transformer block is composed of two sub modules,the Multi Dconv head transposed attention module(MDTA) and the Gated Dconv feed-forward network(GDFN).
9.1 Transformer Block in Res-HAIR Framework
(This section is directly borrowed from the paper of PromptIR. [38]) In this section, we present the block diagram5 of the transformer block and further, elaborate on the details of the transformer block employed in the Res-HAIR framework. The transformer block follows the design and hyper-parameters outlined in [60]
To begin, the input features $\mathbf{X}\in\mathbb{R}^{H_{l}\times W_{l}\times C_{l}}$ are passed through the MDTA module. In this module, the features are initially normalized using Layer normalization. Subsequently, a combination of $1\times 1$ convolutions followed by $3\times 3$ depth-wise convolutions are applied to project the features into Query ($\mathbf{Q}$), Key ($\mathbf{K}$), and Value ($\mathbf{V}$) tensors. An essential characteristic of the MDTA module is its computation of attention across the channel dimensions, rather than the spatial dimensions. This effectively reduces the computational overhead. To achieve this channel-wise attention calculation, the $Q$ and $K$ projections are reshaped from $H_{l}\times W_{l}\times C_{l}$ to $H_{l}W_{l}\times C_{l}$ and $C_{l}\times H_{l}W_{l}$ respectively, before calculating dot-product, hence the resulting transposed attention map with the shape of $C_{l}\times C_{l}$. Bias-free convolutions are utilized within this submodule. Furthermore, attention is computed in a multi-head manner in parallel.
After MDTA Module the features are processed through the GDFN module. In the GDFN module, the input features are expanded by a factor $\gamma$ using $1\times 1$ convolution and they are then passed through $3\times 3$ convolutions. These operations are performed through two parallel paths and the output of one of the paths is activated using GeLU non-linearity. This activated feature map is then combined with the output of the other path using element-wise product.
9.2 Implementation Details of HyperTrans Blocks
This section elaborates on specific details presented in Sec. 3.1.2. Within the Restormer framework [60], each Transformer consists of a total of 12 logical convolutional layers, as depicted in Fig. 5. For code implementation, we utilize 6 convolutional layers to effectively emulate the functionality of 12 convolutional layers within each Transformer Block. Consequently, the generated parameters $\mathbf{w}\in\mathbb{R}^{P}$ described in Sec. 3.1.2 encompass all the parameters necessary for these 6 convolutional layers of a single Transformer Block. The Weights Box $\mathbf{W}\in\mathbb{R}^{N\times P}$, therefore, represents the aggregate parameters for $N$ Transformer Blocks.
Given a fixed Global Information Vector (GIV) $\mathbf{V_{g}}$, it is processed through the Fully-Connected Neural Network (FCNN) followed by a Softmax operation to yield the Selecting Vector $\mathbf{V_{s}}\in\mathbb{R}^{N}$. This vector is subsequently employed to "select" the definitive parameters from the Weights Box. Each Decoder level possesses a distinct Weights Box, which is shared among all the HyperTrans Blocks at that level. It is crucial to highlight that all HyperTrans Blocks across all levels operate using the same GIV derived from a single Classifier. Meanwhile, each single HyperTrans Block is equipped with its own Hyper Selecting Net, resulting in a unique Selecting Vector and, consequently, its own set of parameters for each.
10 Visual Results
In Fig. 6 and Fig. 7 we provide more visual results to show the strong ability of our method for image restoration tasks.
Figure 6: Additional visual results.
Figure 7: Additional visual results